OpenStudy (hang254):

Solve for x. 3/x = x/32

OpenStudy (anonymous):

Cross multiply it and then solve...

OpenStudy (anonymous):

\[x*x = 32*3\] \[x^2 = 96\] Take square root and then get the value of x...

OpenStudy (hang254):

ok

OpenStudy (anonymous):

Able in solving or not??

OpenStudy (hang254):

yes, 9.797

OpenStudy (anonymous):

Right buddy..